OpenStudy (anonymous):

Explain how to determine whether the parabola opens up or down from the quadratic equation.?

OpenStudy (anonymous):

The parabola opens up if the slope is positive The parabola opens down if the slope is negative

OpenStudy (zale101):

if a is positive it opens up if it's negative it opens down y=a(x-h)^2+k

OpenStudy (anonymous):

If the slope is -2x^2, the parabola opens downwards If the slope is 2x^2, the parabola opens up

OpenStudy (zale101):

y=a(x-h)^2+k y=-a(x-h)^2+k